# Measuring Hawking Radiation from Black Hole Morsels in Astrophysical Black Hole Mergers

### v1

@@ Abstract
So far, we have presented simple templates based on BH morsels of the same mass emitted radially as a spherically symmetric shell. An extended mass distribution can modulate and prolong the duration of the TeV photon emission. The geometry of the emission, instead, has a smaller impact on the signal, as BHs emit particles in a spherically symmetric manner. Hence, as long as the morsels are not screened by the merged BH, their Hawking radiation will reach our detectors. Some distortions of the spectrum can be expected if the BH morsels travel with relativistic velocity, in which case their direction of motion with respect to the line of sight becomes important. When BH morsels are emitted in a strong gravitational field, such as for SMBH mergers, their Hawking emission may be affected and a dedicated study is required to quantify such effects, which we leave for future investigations. As a final remark, we recall that BH morsels will most intensely emit at times close to their evaporation and with primary energies exceeding the multi-TeV scale. Besides a correct characterisation of the emission of secondary spectra [66], the Hawking radiation at such energies can be also impacted by the limited validity of the standard model of particle physics, and some new physics effects may induce measurable distortions. The observation of Hawking radiation from BH morsels, therefore, could enlighten us not only about the production of such morsels, but also about particle physics at energies beyond the reach of current and future collider experiments, carrying imprints from new physics based on supersymmetry, composite dynamics, or extra dimensions, to name a few.
In summary, in this letter we have shown that it is possible to observe the Hawking radiation emitted by BH morsels assumed to form in catastrophic astrophysical events such as black hole mergers. GRBs are unique footprints of these asteroid-mass black hole morsels emitted after the merger, with photon energies exceeding the TeV scale. The time delay between the gravitational wave event and the GRB is correlated to the mass distribution of the morsels. We showed that the total morsel mass allowed by the unaccounted merger mass leads to a Hawking induced radiation in photons that is above the sensitivity of atmospheric Cherenkov telescopes such as HESS, LHAASO and HAWC.

### v2

@@ 2.3 Supermassive BH mergers
2.4 Outlook
So far, we have presented simple templates based on BH morsels of the same mass emitted radially as a spherically symmetric shell. An extended mass distribution can modulate and prolong the duration of the TeV photon emission. The geometry of the emission, instead, has a smaller impact on the signal, as BHs emit particles in a spherically symmetric manner. Hence, as long as the morsels are not screened by the merged BH, their Hawking radiation will reach our detectors. Some distortions of the spectrum can be expected if the BH morsels travel with relativistic velocity, in which case their direction of motion with respect to the line of sight becomes important. When BH morsels are emitted in a strong gravitational field, such as for SMBH mergers, their Hawking emission may be affected and a dedicated study is required to quantify such effects, which we leave for future investigations. As a final remark, we recall that BH morsels will most intensely emit at times close to their evaporation and with primary energies exceeding the multi-TeV scale. Besides a correct characterisation of the emission of secondary spectra [71], the Hawking radiation at such energies can be impacted by the limited validity of the standard model of particle physics, and some new physics effects may induce measurable distortions [72, 73]. The observation of Hawking radiation from BH morsels, therefore, could enlighten us not only about the production of such morsels, but also about particle physics at energies beyond the reach of current and future collider experiments, carrying imprints from new physics based on supersymmetry, composite dynamics, or extra dimensions, to name a few.
BH morsels are not the only source of particles being emitted during a merger of compact objects. Several astrophysical mechanisms can produce high-energy transients in the vicinity of BH mergers, including magnetospheric reconnection [74, 75], accretion disk disruption [26, 27], and jet activity [28, 76]. Morsel evaporation events are characterized by a delayed, isotropic, multi-TeV burst coincident with a GW event, lacking persistent afterglow emission or strong beaming. In contrast, disk- or jet-driven GRBs often show beamed emission, multiwavelength afterglows, and may not correlate tightly in time with GW triggers.
3 Conclusions
In this work, we have shown that it is possible to observe the Hawking radiation emitted by BH morsels assumed to form in catastrophic astrophysical events such as astrophysical BH mergers. The resulting GRBs offer unique footprints of these asteroid-mass BH morsels emitted after the merger, with photon energies exceeding the TeV scale. The time delay between the gravitational wave event and the GRB is correlated to the mass distribution of the morsels. We showed that the total morsel mass allowed by the unaccounted merger mass leads to a Hawking induced radiation in photons that is above the sensitivity of atmospheric Cherenkov telescopes such as HESS, LHAASO and HAWC, and that current data already constraints the total mass emitted in morsels to the level of a thousandth solar mass.
@@ Acknowledgements
Appendix A Morsel Creation in Theories beyond GR
Colliding BHs have been identified since long time ago as a source of gravitational waves [77], with the expected outcome of a single merged BH. This expectation follows from a simple application of the area law [78, 79] in GR: while multiple BH final states are thermodynamically allowed, the solution with a single merged one is entropically favoured. However, theories beyond GR have been developed, where fragmentation of black holes occurs. Even within GR, considerations regarding the topology of the horizon typically refer to static solutions of Einstein’s equations [80, 81]: BH mergers are definitely more complicated solutions. Indeed, numerical simulations within GR have already shown the emergence of strong non-linearities in the gravitational field both during the merger [82] and during the successive ringdown phase [83, 84]. Henceforth, one cannot exclude that the extreme environment occurring during the merger could give rise to the production of small BHs, that are successively expelled by the system. Note that non-linearities may also emerge in highly rotating BHs [85, 86, 87, 88].
A.1 Fragmentation of Black Holes
The horizon topology of BHs in 4-dimensional GR is constrained to be spherical [80, 81] and the horizon is resilient to perturbations, rendering the fragmentation of a fully formed horizon unlikely. The situation changes in some extension of general relativity, and we present a non-exhaustive list of possibilities below:
  • In extra dimensional space-time, BH solutions are less restricted. This leads to geometries that exhibit instabilities, even classically, and thus may decay and fragment. For example, it was argued [89, 90] that black strings, i.e. metrics with a horizon geometry of the form $S^{2}\times S^{1}$, and more generally black branes, are unstable under small perturbations. Furthermore, higher dimensional BH solutions are known (e.g. [91]) that allow for arbitrarily high angular momentum per unit mass (thus avoiding the Kerr bound in 4-dimensions [92]). It was argued in [93] that such ultra-spinning BHs exhibit instabilities beyond a certain threshold and thus may also fragment. Such characteristics may survive in models with compact extra dimensions where only gravity propagates [94], hence the products of the horizon fragmentation would resemble BH morsels from a 4-dimensional perspective. In such a scenario, the size of the extra dimensions, which is bound to be below a few microns, would provide a characteristic scale for the morsels.
  • More general higher dimensional configurations have been studied in the context of superstring theory. The spectrum of these BPS solitons (e.g. [95]) generically depends on the position in the moduli space of the theory: across certain co-dimension one surfaces (called walls of marginal stability), single-particle BPS states, resembling (charged) BHs from a 4-dimensional perspective, may decay into more elementary multi-centre configurations. For more information on this wall-crossing phenomenon, we refer the reader to the review article [96] and references therein. We simply remark that string theory generally allows the fragmentation of solitonic solutions, which may be interpreted as the formation of BH morsels.
  • BH horizon instabilities and fragmentation have also been investigated in 4-dimensional modifications of general relativity. For example, in [97] it was shown via entropy arguments that combining gravity extensions of the form f(R) [98] (including monopole dynamics) with the modified uncertainty principle stemming from the existence of a minimum length [99, 100, 101] leads to BH instabilities and fragmentation.
Furthermore, we remark that morsels may also be produced thanks to the presence of matter or fields around the BH horizon. For instance, in the presence of ultra-light bosonic degrees of freedom, such as axion-like dark matter candidates, a Bose condensate forms near the BH horizon via superradiance [102]. It has been shown that, around rotating Kerr BHs, the value of the field can reach values close to the Planck scale [103]. This scenario, therefore, predicts the formation of a shell enveloping the BH, which has been recently considered as a source of accelerated fermions [104]. During a BH merger, the condensate shell may be trapped in between the colliding horizons, be compressed and locally collapse into BH morsels. Note that other effects of superradiance on BH mergers have been studied in [105, 106, 107]. A scenario involving the Higgs field was considered in [17], under the hypothesis that the strong gravitational field near the horizon modifies the Higgs potential and generates a new vacuum configuration. Hence, bubbles of the new vacuum form near the horizon. During a merger, for a short time, the bubbles are trapped in between the merging horizons and collide, hence producing small BHs.
Finally, we would like to mention the effect of a very light scalar field, with Compton wavelength of the same order as the astrophysical BH radius. Such light states may emerge as axion-like particles in many extensions of the standard model of particle physics. It has been shown that a high energy density can accrete in such scalar field around BHs with [103] or without spin [108, 109, 110]. Effectively, this effect would create an envelope around the horizons of the merging BHs at a distance of a few radii and with densities of the order of the BH density [110]. During the merger, the collision of such envelopes could produce morsels, as long as a strong compression is generated during the merger by shock waves of the strong non-linearity in the gravitational field. A detailed study of such mechanism, however, is beyond the scope of this work, and also out of reach from numerical simulation due to the small scales involved. Note on the side that the presence of scalar envelopes could also affect the spectrum of gravitational waves, hence providing a second characteristic signature [105, 106, 107].
We leave a detailed study of the production mechanisms for the morsels to a future publication, as the main focus of the present work is to establish the potential observability of the morsels via their Hawking radiation.
A.2 Back-of-the-Envelop Estimation of Morsel Mass
A different point of view that naively allows for the fragmentation of BHs (particularly in a quantum theoretical setting) stems from the central dogma [18], which states that a BH can be described in terms of an ensemble of quantum states, to construct a toy model leading to the production of morsels. Hence, we model the BH as a bound state of internal components, which follow an appropriate equation of state that balances the gravitational force and prevents further collapse. Qualitatively, this description is fairly similar to that of neutron stars (NSs): it is well known that during NS mergers, a large number of nucleons (the inner components of the NS) are emitted. Numerical simulations estimate that, depending on the conditions of the merger, the total mass ejected in debris can amount up to $10^{-3}$ or $10^{-1}$ solar masses [111]. Stiffer equations of state lead to lesser ejected mass, hence for BH mergers we take conservatively $10^{-3}$ solar masses as a reasonable reference value. To estimate the mass of the ejected BH debris, which we identify with morsels, we need to model the inside of the astrophysical BH. Assuming for simplicity that the mass is uniformly distributed, with density $\rho_{\rm BH}=\frac{3M}{4\pi R^{3}}$, where $R=2GM$ is the BH radius, applying the inner Schwarzschild solution we find that the inner gravitational pressure is constant and equal to
$p_{\rm BH}=-\rho_{\rm BH}=-\frac{3}{8\pi GR}\,.$ (A.1)
Interestingly, this follows the equation of state of vacuum energy. Hence, the interaction among the inner components, which stabilises the BH, would have a positive vacuum energy $V_{\rm int}=-p_{\rm BH}$. From this, we can estimate the typical interaction length:
$l_{\rm int}\sim\frac{1}{V_{\rm int}^{1/4}}=\left(\frac{8\pi}{3}\right)^{1/4}\sqrt{l_{\rm Pl}\ R}=6.8\times 10^{-16}~{}\mbox{m}\times\sqrt{\frac{R}{10~{}\mbox{km}}}\,,$ (A.2)
where $l_{\rm Pl}$ is the Planck length and we used a typical radius of a stellar mass BH. Assuming that the emitted morsels have such a radius, their typical mass would be
$M_{\rm Bm}\sim 4.7\times 10^{11}~{}\mbox{kg}\times\sqrt{\frac{R}{10~{}\mbox{km}}}\,.$ (A.3)
We can consider this as an order of magnitude estimate of the emitted morsels. Clearly, this is a simple minded model, which would need great refinement to properly describe a BH, however it allows us to estimate the key quantities for the study of the morsel phenomenology and observability. A more quantitative study would require numerical simulations of the merger with an incredible precision to resolve the morsels, and with the unknown of the BH equation of state.